can you help me solve 16=2^7x-5
Answer by stanbon(75887)

solve 16=2^7x-5
----
16 = 2^(7x-5)
---
2^4 = 2^(7x-5)
---
7x-5 = 4
7x = 9
x = 9/7
-------------
Cheers,
Stan H.